# GraphToxin: Reconstructing Full Unlearned Graphs from Graph Unlearning

### v1

@@ 2.5 Problem Statement
We formally define the full graph reconstruction attack against graph unlearning (GU-FGRA) as follows. Specifically, for a single node removal, attackers aim to identify the optimal attack model $\mathcal{F}_{{att}_{\Delta\mathcal{G}_{v}}}={}^{*}\mathcal{F}^{-1}_{\Delta\mathcal{G}_{v}}(\Delta\nabla\mathcal{L}(\theta_{\Delta\mathcal{G}_{v}}))$ to recover the full unlearned graph $\mathcal{G}_{{d}_{v}}$, where $\Delta\nabla\mathcal{L}(\theta_{\Delta\mathcal{G}_{v}})$ denotes the difference between the gradients of the original and unlearned models ($\nabla\mathcal{L}(\theta_{ori})$ and $\nabla\mathcal{L}(\theta_{{un}_{\Delta_{\mathcal{G}_{v}}}})$), and $\mathcal{L}$ represents the empirical risk function. $\mathcal{G}_{d_{v}}=(\mathcal{V}_{{rec}_{v}},\mathcal{E}_{\mathcal{N}^{1}_{v}},\mathcal{X}_{{rec}_{v}})$, where recovery nodes $\mathcal{V}_{{rec}_{v}}=[v,\mathcal{N}^{1}_{v}]$ contain the deleted node $v$ and its affected 1-hop neighbors $\mathcal{N}^{1}_{v}$. Please note that $\mathcal{G}_{d_{v}}\neq\Delta\mathcal{G}_{v}$ since there is no isolated node in $\mathcal{G}$, $|\mathcal{G}_{d_{v}}|>|\Delta\mathcal{G}_{v}|$. Moreover, recovering $\Delta\mathcal{G}_{v}$ can be simply treated as a combination of $\mathcal{D}$ times attribute inference attacks of a single node $v$ and a graph topology reconstruction attack of its associated links $\mathcal{N}^{1}_{v}$ (i.e., a star graph). However, recovering $\mathcal{G}_{d_{v}}$ is significantly more complicated, as it further requires the simultaneous full feature recovery of multiple nodes. For multiple node removals, attackers strive to reveal $\mathcal{G}_{d}$ through the optimal attack model $\mathcal{F}_{{att}_{\Delta\mathcal{G}}}$. Similarly, $\mathcal{G}_{d}=(\mathcal{V}_{{rec}_{\mathcal{V}_{d}}},\mathcal{E}_{\mathcal{N}^{1}_{\mathcal{V}_{d}}},\mathcal{X}_{{rec}_{\mathcal{V}_{d}}})\neq\Delta\mathcal{G}$, where $\mathcal{V}_{d}$ is the node deletion set. For black-box GU-FGRA, attackers exploit the posterior probabilities $P=(p_{1},p_{2},\dots,p_{|Y|})$ obtained before and after node removals to estimate the gradient difference $\Delta\hat{\nabla}\mathcal{L}(\theta_{\Delta\mathcal{G}_{v}})$ or $\Delta\hat{\nabla}\mathcal{L}(\theta_{\Delta\mathcal{G}})$, and then adopt the identical procedure used in the white-box setting.
@@ 3.1.1 Attacker’s Goal
As mentioned above, attackers aim to fully reconstruct the unlearned graph—including the removed nodes, their affected nodes, their topological structures, and node features—rather than a simple combination of multiple feature recovery and neighboring link stealing attacks either for a single node or multiple node removals.
@@ 3.1.2 Attacker’s Knowledge and Capabilities
Graph unlearning generally involves three parties in practice: the model owner, the server and the users. The model owner trains a GNN on a private graph and then uploads it to the server. The server then deploys the model and packs it into an API for users to use, and the users can flexibly send queries to the server and receive prediction results [first_unlearn_attack, unlearn_membership_attack]. In compliance with legal regulations, upon receiving an unlearning request from a participant (i.e., a specific node in the training graph), the model owner must adopt graph unlearning methods to unlearn any information associated with that participant, and then upload the unlearned model and delete the previous version. However, attackers can exploit the time gap between model updates to save a copy of the deleted model or perform model extraction attacks to steal it. Therefore, it is realistic to assume that attackers have access to both the original and unlearned models [unlearn_membership_attack, first_unlearn_attack]. We also assume attackers can master the number of changed nodes, edges, and corresponding labels. This is practical since such information can be obtained through simple count queries or by inspecting publicly released technical documents. These documents—which often detail data statistics and model information (e.g., node and edge size, label distribution, model name)—are common on MLaaS platforms to facilitate proper model use. Moreover, attackers can employ existing label inference techniques to recover labels, which serve as a cornerstone for subsequent reconstruction attacks [first_unlearn_attack, label_inf].
In contrast to existing graph reconstruction attacks, our threat model strictly assumes no access to partial node features, graph topology, or shadow datasets. This assumption is realistic, as private training graphs are proprietary and constitute valuable assets of companies and institutions, necessitating strong protections against potential attackers.
In reality, attackers can be an internal actor, such as honest-but-curious server administrators, they faithfully fulfill their responsibilities to serve users but are curious about the unlearned content upon new unlearning requests [first_unlearn_attack]. They normally deploy the unlearned model but secretly archive the old version. Alternatively, attackers can also be malicious external users, they can conduct model extraction attacks to obtain two copycats of the original and unlearned models. Even though these substitutes can leak much more information, we strictly assume that attackers only have access to the gradients of the original and unlearned models, without prior knowledge of the unlearning method or any partial information of the deleted content. Furthermore, auditors can act as a red team to launch GU-FGRA with only the minimal required knowledge in a data-free setting, enabling them to discover vulnerabilities and anomalies to improve the security and privacy of AI systems.
Under the black-box setting, we assume that attackers can only obtain the posterior probabilities of the original and unlearned GNNs. This more realistic assumption significantly increases the attack range in practice, as any attacker can impersonate as a normal user to query the models and exploit these posteriors to launch GU-FGRA.
@@ 3.2 Attack Framework Design
We introduce GraphToxin, the first full graph reconstruction attack against graph unlearning. The framework overview is shown in Figure 1. GraphToxin contains three modules: gradient matching, curvature matching and feature smoothness, detailed as follows.
@@ 3.2.1 Gradient Matching
Gradient matching is the foundational component of gradient inversion attacks tailored for federated learning. Given the model gradients for a single data point, gradient matching iteratively optimize the dummy input by minimizing the distance between its dummy gradients and the actual model gradients. The optimized dummy input finally “matches” the ground-truth data point. This matching degree is typically quantified using cosine similarity or L2 distance.
In the graph domain, gradient matching denotes the process of reducing the discrepancy between the dummy gradients of a dummy graph and the ground-truth model gradients, it can be formally defined as:
$\tilde{\mathcal{G}}=\arg\min\limits_{\tilde{\mathcal{G}}}\mathrm{D}(\nabla\mathcal{L}(\mathcal{F}_{\theta}(\tilde{\mathcal{G}}),\mathcal{Y}),\nabla\mathcal{L}(\mathcal{F}_{\theta}(\mathcal{G}),\mathcal{Y}))$ (2)
For simplicity, we denote $\mathcal{L}(\mathcal{F}_{\theta}(\mathcal{G}),\mathcal{Y})$ as $\mathcal{L}(\theta)$, the gradients as $\nabla\mathcal{L}(\theta)$ and the approximated gradients as $\tilde{\nabla}\mathcal{L}(\theta)$.
Since graph unlearning requires removing the deleted node $v$’s influence, the difference between the gradients of the original model $\nabla_{ori}\mathcal{L}(\theta)$ and the unlearned model $\nabla_{un}\mathcal{L}(\theta^{\ast})$ consequently encodes the unlearned information $X_{v}$ and $\mathcal{N}^{1}_{v}$. Therefore, we exploit this difference $\nabla_{ori}\mathcal{L}(\theta)-\nabla_{un}\mathcal{L}(\theta^{\ast})$ ($\Delta\mathcal{L}(\mathcal{G}_{d},\theta,\theta^{\ast})$ hereinafter) to recover the full unlearned graph $\mathcal{G}_{d}$, which can be represented as:
$\tilde{\mathcal{G}_{d}}=\arg\min\limits_{\tilde{\mathcal{G}}_{d}}\mathrm{D}(\tilde{\Delta}\mathcal{L}(\tilde{\mathcal{G}_{d}},\theta,\theta^{\ast}),\Delta\mathcal{L}(\mathcal{G}_{d},\theta,\theta^{\ast}))$ (3)
@@
For the choice of distance function $\mathrm{D}$, existing gradient inversion attacks commonly use cosine similarity [first_gradient, first_unlearn_attack] or L2 distance [deep_leakage_gradients]. One empirical investigation [l2_cosine_insight_gradient] favors L2 distance for its superior attack performance while another [fastconvergentgradient] shows that cosine distance correlates more strongly with the reconstruction quality. Inspired by Fan et al. [realistic_gradient], we combine both metrics to leverage their complementary strengths. Specifically, we adopt Mean Squared Error (MSE), an L2-based metric, to measure the absolute distance in the high-dimensional gradient space while mitigating the magnitude influences of the huge number of gradients, and leverage cosine similarity for direction alignment. Following Fan et al. [realistic_gradient], we treat these two losses equally.
@@
$\begin{split}\mathfrak{L}_{grad}&=1-\frac{\langle\tilde{\Delta}\mathcal{L}(\tilde{\mathcal{G}_{d}},\theta,\theta^{\ast}),\Delta\mathcal{L}(\mathcal{G}_{d},\theta,\theta^{\ast})\rangle}{\|\tilde{\Delta}\mathcal{L}(\tilde{\mathcal{G}_{d}},\theta,\theta^{\ast})\|\cdot\|\Delta\mathcal{L}(\mathcal{G}_{d},\theta,\theta^{\ast})\|}\\ &+\frac{1}{M}\sum\limits_{j=1}^{M}(\tilde{\Delta}\mathcal{L}(\tilde{\mathcal{G}_{d}},\theta,\theta^{\ast})_{j}-\Delta\mathcal{L}(\mathcal{G}_{d},\theta,\theta^{\ast})_{j})^{2}\end{split}$ (4)
@@
However, gradient matching is necessary yet insufficient for recovery. The corresponding analysis is detailed below.
Necessity Gradients manifest the learning behaviors of a trained model and encode high-dimensional, per-sample fingerprints of the training graph. Gradient matching exploits these fingerprints to recover the full graph by synthesizing a candidate graph and iteratively adjusting it until its generated gradients match the observed ones. This process effectively mimics the model’s learning behaviors and “back-propagates” information from the gradient space to the input space. Moreover, since gradients quantify how sensitive the model’s loss is to parameter changes, gradient matching inherently seeks to a candidate graph that elicits the same sensitivities in the trained model and implicitly prioritizes the features and structures that are most salient to the model’s learning process and ultimately contribute most to its decision boundary.
Insufficiency Gradient matching only provides coarse proxy signals, as multiple different candidate graphs can yield very similar gradients, hence, attackers may match gradients but still recover wrong node features or graph topology. This ambiguity stems from inherent GNN properties: since GNNs are permutation-invariant, isomorphic graphs can produce identical information flow and hence identical gradients. Moreover, complex graph topology and node features are entangled during GNN message passing, severely compromising the identifiability of individual graph components within a single gradient signal. These issues are particularly magnified in graph unlearning. First, attackers exploit the gradient difference between the original and unlearned GNNs for full graph recovery, which often has a much smaller magnitude than pure gradients and thus providing more ambiguous guidance for disentangling the attribution of each graph component. Second, graph unlearning methods can introduce noises through regularization or mixed optimization objectives, hereby corrupting the unlearning signal. Most importantly, GU-FGRA aims not only to recover $X_{v}$ and $\mathcal{E}_{\mathcal{N}^{1}_{v}}$, but also to reveal the node features $X_{\mathcal{N}_{v}^{1}}$ of the affected nodes. $\Delta\mathcal{L}(\mathcal{G}_{d},\theta,\theta^{\ast})$ can explicitly reflect $X_{v}$ and $\mathcal{E}_{\mathcal{N}^{1}_{v}}$, but not $X_{\mathcal{N}_{v}^{1}}$. The precise recovery of $X_{\mathcal{N}_{v}^{1}}$ for a single node removal still poses extra challenges, not to mention the extension to multiple node removals.
@@ 3.2.2 Curvature Matching
Through the preceding analysis, we identify the key limitation in standard gradient matching: solely relying on the gradient difference matching before and after graph unlearning can be ambiguous. As depicted in the illustrative example in Figure 2, we simulate the unlearning effect (reflected in $\mathcal{L}_{2}$) by adding noises and slightly changing the polynomial coefficients of a convex loss function $\mathcal{L}_{1}$. As shown in Figure 2.A, the gradients of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are almost identical and their small gradient difference is reflected in Figure 2.B. However, we identify two distinct points (marked in blue and pink) that produce identical gradient differences, yet exhibit the substantially different curvatures of the loss landscape as shown in Figure 2.C and D. To overcome this ambiguity and provide fine-grained signals for full unlearned graph recovery, we introduce a novel curvature matching module.
Figure 2: Illustrative Example of the Ambiguity of Gradient Difference Matching
Assumption 1 (Empirical Risk Function).
We assume that $\mathcal{L}(\cdot)$ is twice continuously differentiable and the Hessian $H=\nabla^{2}_{ori}\mathcal{L}(\theta)$ is invertible and symmetric w.r.t. parameter $\theta$. Following previous studies [hessian, hessian_influence], we add a damping term $\lambda I$ to $H$ if $\mathcal{L}(\cdot)$ is non-convex.
Assumption 2 (Small Parameter Change).
We empirically find that this change is small (reflected in subsequent experiments). This finding aligns with theoretical works proving the change to be bounded [GIF, chien2022certified].
When the above assumptions hold, the first-order Taylor expansion of the gradient function $\nabla\mathcal{L}(\cdot)$ around a stationary point $\theta$ is:
$\begin{split}\nabla_{un}\mathcal{L}(\theta^{\ast})\approx\nabla_{ori}\mathcal{L}(\theta)+\nabla^{2}_{ori}\mathcal{L}(\theta)(\theta^{\ast}-\theta)\end{split}$ (5)
GU-FGRA inherently generates a full recovered graph whose gradients can explain the parameter change [first_unlearn_attack]. Hence:
$\Delta\tilde{\theta}\approx\theta^{\ast}-\theta$ (6)
where $\Delta\tilde{\theta}$ is the parameter induced by the dummy graph. Since the parameter $\theta$ of the original model is fixed, w.l.o.g., for GU-FGRA, the attackers aim to find the parameter $\tilde{\theta}$ so that $\Delta\tilde{\theta}=\tilde{\theta}-\theta$.
By combining Equation (5) and Equation (6), we state the following theorem:
Theorem 1 (Fine-grained Curvature Matching).
When the above assumptions hold, matching the ground-truth gradient difference $\Delta\mathcal{L}(\mathcal{G}_{d},\theta,\theta^{\ast})$ is equivalent to minimizing $(\Delta_{syn}-\Delta_{obs})^{T}H^{-1}(\Delta_{syn}-\Delta_{obs})$. where $\Delta_{syn}=\tilde{\Delta}\mathcal{L}(\tilde{\mathcal{G}_{d}},\theta,\theta^{\ast})$, $\Delta_{obs}=\Delta\mathcal{L}(\mathcal{G}_{d},\theta,\theta^{\ast})$.
Proof. Based on Equation (5), we get $\Delta_{obs}\approx H(\theta^{\ast}-\theta)$. Since H is invertible, $\theta^{\ast}-\theta\approx H^{-1}\Delta_{obs}$. As for Equation (6), under the same linearized relation in Equation (5), $\tilde{\theta}-\theta\approx H^{-1}\Delta_{syn}$. Therefore, the attack goal is transformed to minimize the difference between $H^{-1}\Delta_{obs}$ and $H^{-1}\Delta_{syn}$. Formally, $H^{-1}(\Delta_{syn}-\Delta_{obs})\approx 0$. Take the $H^{1/2}$ to the left-hand and right-hand side, then we get $H^{-1/2}(\Delta_{syn}-\Delta_{obs})\approx 0$. Consider the squared L2 norm, the previous approximation equals to $(\Delta_{syn}-\Delta_{obs})^{T}H^{-1}(\Delta_{syn}-\Delta_{obs})$, hereby completing the proof.
@@
However, utilizing $H$ is computationally infeasible and expensive as it requires calculating and storing second-order derivatives. We leverage the Fisher Information Matrix $F$ as the surrogate. Specifically, $F$ is defined as the covariance matrix of the gradient of the log-likelihood function with respect to the parameters $\theta$, quantifying the sensitivity of the model’s predicted probability distribution to small changes in $\theta$. According to Rame et al. [fisher], $F$ approximates $H$ with probably bounded errors under regular conditions.
In practice, since the full $F$ involves an expectation over all possible outputs, the empirical Fisher Information $\tilde{F}$ is typically utilized as an efficient approximation [empirical_fisher]. Formally, $\tilde{F}$ is denoted as:
$\tilde{F}=\frac{1}{N_{d}}\sum\limits_{i=1}^{N_{d}}(\nabla\log p(y_{i}|\tilde{\mathcal{G}}_{d},\theta))(\nabla\log p(y_{i}|\tilde{\mathcal{G}}_{d},\theta))^{T}$ (7)
where $N_{d}=|\mathcal{V}_{{rec}_{\mathcal{V}_{d}}}|$ is the node size of $\mathcal{G}_{d}$.
@@
Replacing $H$ with $\tilde{F}$ in $(\Delta_{syn}-\Delta_{obs})^{T}H^{-1}(\Delta_{syn}-\Delta_{obs})$, we derive the following curvature matching loss:
$\mathfrak{L}_{curv}=(\Delta_{syn}-\Delta_{obs})^{T}\tilde{F}^{-1}(\Delta_{syn}-\Delta_{obs})$ (8)
The loss $\mathfrak{L}_{curv}$ naturally encodes the curvature of the loss landscape, providing much more precise guidance for full unlearned graph recovery. Not all gradient discrepancies between the synthetic and the full unlearned graphs are equally informative; some may stem from random noises while others are highly indicative of the ground-truth features and underlying graph structures. Introducing $\mathfrak{L}_{curv}$ can penalize the directions of high curvature, thereby biasing the optimization process toward plausible solutions with smooth landscapes and avoiding overfitting to noisy gradient directions. Moreover, the inclusion of $F$ can capture the global features and structural dependencies and reflect how they are propagated during GNN message-passing and consequently projected into posterior probabilities. Compared to the simplex gradient matching, curvature matching encompasses a more holistic information flow, therefore yielding more realistic and accurate recovery.

### v2

@@ 2.5 Problem Statement
We formally define the full graph reconstruction attack against graph unlearning (GU-FGRA). Specifically, for single-node removal, a GU-FGRA aims to recover the full unlearned graph $\mathcal{G}_{{u}_{v}}=(\mathcal{V}_{{u}_{v}},\mathcal{E}_{u_{v}},X_{{u}_{v}})$, where the recovered nodes $\mathcal{V}_{{u}_{v}}=[v,\mathcal{N}^{1}_{v}]$ contain the removed node $v$ and its affected 1-hop neighbors $\mathcal{N}^{1}_{v}$. Please note that $\mathcal{G}_{u_{v}}\neq\Delta\mathcal{G}_{v}$. Unlike $\Delta\mathcal{G}_{v}$, $\mathcal{G}_{u_{v}}$ additionally recovers node features of its 1-hop neighbors. Moreover, recovering $\Delta\mathcal{G}_{v}$ can be simply treated as a combination of $d$ times attribute inference attacks of a single node $v$ and a graph topology reconstruction attack of its associated links $\mathcal{N}^{1}_{v}$, where $d$ denotes the number of node features. In contrast, recovering $\mathcal{G}_{u_{v}}$ is significantly more complicated, as it further requires the simultaneous full feature recovery of multiple recovered nodes. For multiple-node removal, attackers strive to reconstruct $\mathcal{G}_{u}=(\mathcal{V}_{u},\mathcal{E}_{u},X_{u})$. Similarly, $\mathcal{G}_{u}\neq\Delta\mathcal{G}$ where $\mathcal{V}_{u}=[\mathcal{V}_{\mathcal{M}},\mathcal{N}^{1}_{\mathcal{M}}]$ contains multiple removed nodes $\mathcal{M}$ and their 1-hop neighbors $\mathcal{N}^{1}_{\mathcal{M}}$. Under white-box settings, the adversary leverages model gradients $\nabla\mathcal{L}$ to recover the full unlearned graph for both single-node or multiple-node removal, while under black-box settings, she/he exploits the posterior probabilities $P=(p_{1},p_{2},\dots,p_{|Y|})$ obtained before and after node removals to achieve the same objective.
@@ 3 Attack Framework Design
In this section, we formalize the threat model and introduce GraphToxin, the first data-free full graph reconstruction attack against graph unlearning, followed by an evaluation framework to systematically assess it.
@@ 3.1.1 Attacker’s Goal
The attacker aims to fully reconstruct the unlearned graph, including the removed nodes, their affected neighborhoods, the induced topological structures, and associated node features.
@@ 3.1.2 Attacker’s Knowledge and Capabilities
We consider a standard MLaaS deployment involving a model owner, a server, and users. The model owner trains a GNN on a private graph and uploads the GNN to the server. The server then deploys the model and packs it into an API for users to use [first_unlearn_attack, unlearn_membership_attack]. Upon receiving an unlearning request from a participant (i.e., a specific node in the private training graph), the model owner adopts a graph unlearning method to remove any information associated with that participant (i.e., its node features and neighboring links), releases the unlearned GNN, and deletes the previous version. However, the involvement of multiple parties and time gaps between model updates expand the attack surface, where the adversary can exploit to reconstruct the full unlearned graphs in a data-free manner.
Our attack setting fundamentally differs from existing reconstruction attacks. They assume that the adversary either has partial access to the original training graphs or unlearned graphs, or can obtain shadow graphs. However, these assumptions are unrealistic: first, original training graphs are proprietary and constitute valuable assets of companies and institutions, necessitating strong protections against potential attackers; second, shadow graphs typically demand same or similar distributions from the original training graphs. Since training graphs are usually protected, it is hard for the adversary to acquire the distribution information. Therefore, GU-FGRA should be strictly operated under a data-free setting.
We consider two types of attackers: the honest-but-curious server and malicious user. Following Hu et al., [first_unlearn_attack], the server has the white-box access to the original and unlearned GNNs, but only restricted to model gradients, without any prior knowledge of unlearning methods, GNN architecture, and training or unlearned graph distributions. This assumption is realistic as the server can save copies of all uploaded models or archive the deleted versions for legitimate reasons, including but not limited to backup, auditing, debugging, or rollback after accidental deletion. Meanwhile, the malicious user only has the black-box access to the original and unlearned GNNs. She/he can submit model queries to the server, and the server will faithfully respond these queries and send posterior-based model predictions back. The user can conduct data-free model stealing attacks to construct surrogate models based on collected query-response pairs [data_free_msa_gnn]. In addition, auditors can also act as a red team to launch GU-FGRA with only the minimal required knowledge in a data-free setting, enabling them to discover vulnerabilities and anomalies to improve the security and privacy of graph-based AI systems.
We assume that the adversary has knowledge of the number of removed nodes and edges, as well as changes in label distributions. This assumption is practical since such information can be exposed through simple count queries, system logs or publicly available technical documents, which are commonly provided on MLaaS platforms to facilitate proper model usage. Moreover, attackers can employ existing label inference techniques to recover labels, which serve as a cornerstone for subsequent reconstruction attacks [first_unlearn_attack, label_inf].
@@ 3.2 Attack Framework Design
We introduce GraphToxin and present its framework overview in Figure 1. GraphToxin contains three modules: Gradient Matching to extract holistic information for the full unlearned graph recovery, Curvature Matching to provide more fine-gradient recovery signals and Feature Smoothness to capture and enforce similarity among connected nodes within the unlearned graph. We detail each module as follows.
@@ 3.2.1 Gradient Matching
Design Rationale Gradient matching is the foundation of gradient inversion attacks mainly tailored for federated learning. Given the gradients with respect to model parameters, gradient matching first initializes a dummy input and then iteratively optimizes it by minimizing the distance between the gradients induced by the dummy input and the observed model gradients. Since our attack settings are strictly data-free and only assume the adversary access to actual or estimated gradients of the original and unlearned models, gradient matching serves as a natural starting point.
In the graph domain, gradient matching searches for a candidate graph whose induced gradients are consistent with the observed ones. Rather than assuming a direct correspondence between gradients and individual graph components, this process imposes a constraint that narrows the space of plausible graph reconstructions. However, gradient matching cannot be trivially applied to GU-FGRA, as the attack involves two gradient sets from the original and unlearned models. Naively applying gradient matching to each gradient set introduces redundant optimization paths while discarding the unlearning-induced discrepancy that carries the most informative recovery signals. Due to the ambiguity of gradient-to-graph mappings, independently reconstructed graphs are not guaranteed to align, hindering the reliable identification of unlearned graph components.
These limitations motivate a discrepancy-driven design that directly operates on gradient differences between the original and unlearned GNNs, rather than on separate gradient sets.
Gradient Matching in GU-FGRA For clarity, we begin with the single-node removal setting and discuss the extension to multiple-node removal in a separate section.
Formally, let $\nabla\mathcal{L}$ and $\nabla\mathcal{L}^{-v}$ denote the gradients of the original and unlearned GNNs, respectively, where v-v represents an unlearning request on a single node $v$. We define the observed gradient difference as $\Delta\nabla\mathcal{L}=\nabla\mathcal{L}-\nabla\mathcal{L}^{-v}$, which captures unlearning-induced signals associated with the removed node $v$ and its affected neighborhoods $\mathcal{N}^{1}_{v}$. The adversary first initializes a dummy graph with the fixed unlearning graph size, and then optimizes it by minimizing the distance between the dummy and observed gradient difference to recover the full unlearned graph $\mathcal{G}_{u_{v}}=(\mathcal{V}_{u_{v}},\mathcal{E}_{u_{v}},\mathcal{X}_{u_{v}})$, where $\mathcal{V}_{u_{v}}=[v,\mathcal{N}^{1}_{v}]$. The attack objective is represented as:
$\tilde{\mathcal{G}}_{u_{v}}=\arg\min\limits_{\tilde{\mathcal{G}}_{u_{v}}}\mathcal{D}(\Delta\tilde{\nabla}\mathcal{L},\Delta\nabla\mathcal{L})$ (2)
@@
where $\Delta\tilde{\nabla}\mathcal{L}$ denotes the gradients induced by the dummy graph. For the choice of the distance function $\mathcal{D}$, existing gradient matching commonly uses cosine similarity [first_gradient, first_unlearn_attack] or L2 distance [deep_leakage_gradients]. Some prior studies favor L2 distance for its superior attack performance while others [fastconvergentgradient] show that cosine similarity correlates more strongly with reconstruction quality. Inspired by Fan et al. [realistic_gradient], we combine both metrics to leverage their complementary strengths. Specifically, we adopt Mean Squared Error (MSE), a normalized form of L2 distance, to penalize the absolute deviations and mitigate the influence of magnitude amplification in the high-dimensional gradient space. Meanwhile, cosine similarity is used for direction consistency and semantic alignment. Following Fan et al. [realistic_gradient], we treat these two losses equally:
@@
$\begin{split}\mathfrak{L}_{grad}&=1-\frac{\langle\Delta\tilde{\nabla}\mathcal{L},\Delta\nabla\mathcal{L}\rangle}{\|\Delta\tilde{\nabla}\mathcal{L}\|\cdot\|\Delta\nabla\mathcal{L}\|}+\frac{1}{M}\sum\limits_{i=1}^{M}\left((\Delta\tilde{\nabla}\mathcal{L})_{i}-(\Delta\nabla\mathcal{L})_{i}\right)^{2}\end{split}$ (3)
@@ 3.2.2 Curvature Matching
Motivation Gradient matching provides holistic information for the full unlearned graph recovery, however, it is not sufficient to extract more fine-grained recovery signals. The following illustrative example demonstrates how gradient matching leads to ambiguity in GU-FGRA and motivate the design of curvature matching.
Figure 2: Example of the Ambiguity of Gradient Matching
@@
As depicted in Figure 2, we consider a convex loss function $\mathcal{L}_{1}$ and obtain $\mathcal{L}_{2}$ by adding noises and slightly changing the polynomial coefficients of $\mathcal{L}_{1}$ to simulate the effect of unlearning. As shown in Fig.2.A and B, the two losses induce nearly identical gradients and small gradient difference. However, we identify two distinct points (marked in blue and pink) that produce the identical gradient difference, yet exhibit the substantially different curvatures of the loss landscape as shown in Fig.2.C and D. This example highlights a fundamental limitation of gradient matching: multiple candidates can satisfy the same gradient difference constraint. In graph learning, this ambiguity is further amplified by the permutation invariance of GNNs, under which different graph structures can yield identical information propagation and gradients. Moreover, GU-FGRA aims to recover not only the removed node and its incident edges, but also the features of affected neighboring nodes. While gradient difference can reflect the removal of a node and local connectivity changes, they impose insufficient constraints to uniquely characterize neighboring node features. To overcome this ambiguity and provide fine-grained signals for full unlearned graph recovery, we next introduce a novel curvature matching module.
Curvature Matching in GU-FGRA We formalize curvature matching as a curvature-aware constraint operating on gradient difference. Detailed theorem assumptions and proof are provided in Appendix A.
Theorem 1 (Curvature-aware difference Matching).
Under standard smoothness and local approximation assumptions, minimizing the gradient difference reduces to minimizing the curvature-weighted difference $(\Delta\tilde{\nabla}\mathcal{L}-\Delta\nabla\mathcal{L})^{T}H^{-1}(\Delta\tilde{\nabla}\mathcal{L}-\Delta\nabla\mathcal{L})$.
@@
In practice, the Hessian matrix $H$ is computationally infeasible as it requires calculating and storing second-order derivatives. We leverage the Fisher Information Matrix $F$ as the surrogate. Specifically, $F$ is defined as the covariance matrix of the gradient of the log-likelihood function with respect to model parameters $\theta$, quantifying the sensitivity of the model’s predicted probability distribution to small changes in $\theta$. According to Rame et al. [fisher], $F$ approximates $H$ with probably bounded errors under regular conditions. However, since the full $F$ involves an expectation over all possible outputs, the empirical Fisher Information $\tilde{F}$ is typically utilized as an efficient approximation [empirical_fisher]. Formally, $\tilde{F}$ is denoted as:
$\tilde{F}=\frac{1}{N_{u}}\sum\limits_{i=1}^{N_{u}}(\nabla\log p(Y_{i}|\tilde{\mathcal{G}}_{u},\theta))(\nabla\log p(Y_{i}|\tilde{\mathcal{G}}_{u},\theta))^{T}$ (4)
where $N_{u}=|\mathcal{V}_{u}|$ is the node size of $\mathcal{G}_{u}$.
@@
We then replace $H$ with $\tilde{F}$ in $(\Delta\tilde{\nabla}\mathcal{L}-\Delta\nabla\mathcal{L})^{T}H^{-1}(\Delta\tilde{\nabla}\mathcal{L}-\Delta\nabla\mathcal{L})$ and derive the following curvature matching loss:
$\mathfrak{L}_{curv}=(\Delta\tilde{\nabla}\mathcal{L}-\Delta\nabla\mathcal{L})^{T}\tilde{F}^{-1}(\Delta\tilde{\nabla}\mathcal{L}-\Delta\nabla\mathcal{L})$ (5)
$\mathfrak{L}_{curv}$ naturally encodes the curvature of the loss landscape, providing more fine-grained signals for full unlearned graph recovery. Not all gradient difference are equally informative: some may stem from noises or redundant parameterization, while others are highly indicative of the ground-truth features and underlying graph structures. By incorporating $\tilde{F}$, $\mathfrak{L}_{curv}$ downweights the directions of high curvature, thereby biasing the optimization toward smoother and more stable directions in parameter space. Furthermore, $\tilde{F}$ is defined with respect to model predictions and thus implicitly reflects how node features and graph structures are propagated through message passing and projected into posterior probabilities. This complementary perspective provides sensitivity-aware guidance that better aligns with the model’s local behavior.